# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v3

@@ 3.5 Optimization Algorithm
The FairPO framework solves the minimax objective (Eq. 18) iteratively. A high-level overview is in Algorithm 1, with full details in Algorithm 2 (see Appendix B). In each training iteration, an instance and a label $r$ are sampled. The corresponding loss ($\mathcal{L}_{\mathcal{P}}^{(s)}$ or $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ via Eq. 11 or Eq. 17) and its gradient are computed based on whether $r$ is in the privileged set $\mathcal{P}$ or non-privileged set $\bar{\mathcal{P}}$.
The group weights $\alpha^{(s+1)}$ are updated using a mirror ascent step. A crucial aspect for the stability and effectiveness of this step is loss scaling. Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version. Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted $\bar{\mathcal{L}}_{g}^{\text{avg}}$. The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability. This normalization prevents instability due to scale disparities in group loss values and makes the optimization more sensitive to relative improvements or regressions, rather than absolute magnitudes.
This scaled loss, say $\Delta\tilde{\mathcal{L}}_{g}^{(s)}$, is then used in the exponential update: $\alpha_{g}^{\text{new}}\leftarrow\alpha_{g}^{(s)}\exp(\eta_{\alpha}\Delta\tilde{\mathcal{L}}_{g}^{(s)})$, followed by normalization so $\sum_{g}\alpha_{g}^{\text{new}}=1$. This scaling normalizes loss magnitudes across groups, makes weight updates sensitive to significant performance changes relative to recent history, and improves overall stability of the $\alpha$ dynamics. Subsequently, model parameters $\mathbf{w}^{(s+1)}$ are updated via mirror descent, using a gradient that combines group gradients weighted by the adaptively balanced $\alpha^{(s+1)}$.
@@
6:     Compute privileged loss $\mathcal{L}_{\mathcal{P}}^{(s)}$ for $(x_{i},r)$ (Eq. 11, conditionally using preference loss like Eq. 8/9 or BCE Eq. 10).
@@
13:Full details are in Algorithm 2 (see Appendix B).
@@ 4 Experimental Setup
We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014 (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009). For our fairness setup, we define the privileged group ($\mathcal{P}$) as the 20% least frequent labels in each training set, with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$). Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021), where we fine-tune label-specific classifier heads on top of its features. The reference model parameters $\hat{\mathbf{w}}$ are obtained from a standard supervised fine-tuning (SFT) of this architecture with a Binary Cross-Entropy (BCE) loss.
Performance is assessed separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ sets using standard metrics: Mean Average Precision (mAP), Sample F1, Exact Match Ratio (EMR), and Accuracy. We compare FairPO against three baselines: (1) BCE-SFT, which also serves as our reference model; (2) BCE-SFT + Re-Weighting (RW), a simple loss up-weighting for $\mathcal{P}$ labels; and (3) Group DRO + BCE (Sagawa et al., 2020), which minimizes the worst-group loss. We test three variants of our framework, differing in their preference loss for privileged labels: FairPO-DPO (Eq. 11), FairPO-SimPO (Eq. 12), and FairPO-CPO (Eq. 14). Comprehensive details on datasets, preprocessing, model architecture, baseline implementations, and hyperparameter tuning are deferred to Appendices C, D, and E.
@@ 5 Results and Analysis
Table 1: Performance comparison on MS-COCO. $\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%). Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
@@
Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 86.32 90.65 61.43 64.89 94.89 98.12 35.78 36.98 Ref BCE SFT + RW 87.25 89.85 62.68 64.11 95.95 97.93 47.43 33.81 +0.93 GDRO + BCE 87.92 90.41 62.31 64.75 95.72 98.05 46.12 35.16 +1.60 FairPO-DPO 88.02 89.97 63.45 63.65 97.89 98.04 62.19 35.12 +1.70 FairPO-SimPO 87.67 88.76 62.34 63.12 95.69 97.78 45.32 32.34 +1.35 FairPO-CPO 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44
Table 2: Performance comparison on NUS-WIDE. $\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%). Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
@@
Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 63.53 70.24 48.12 55.83 91.51 95.22 19.32 11.56 Ref BCE SFT + RW 65.12 69.14 49.51 54.73 92.33 94.81 21.23 10.32 +1.59 GDRO + BCE 64.84 69.91 49.13 55.62 92.11 95.13 21.02 11.34 +1.31 FairPO-DPO 66.34 69.05 51.71 54.52 93.92 95.04 27.91 11.21 +2.81 FairPO-SimPO 64.11 68.03 48.82 53.81 91.94 94.52 20.18 10.19 +0.58 FairPO-CPO 67.12 69.83 52.21 55.24 94.31 95.12 31.87 11.25 +3.59
Tables 1 and 2 detail our findings. The standard BCE-SFT baseline confirms the fairness problem central to our motivation: on MS-COCO, a significant performance gap exists between the privileged ($\mathcal{P}$) labels (86.32% mAP) and the non-privileged ($\bar{\mathcal{P}}$) labels (90.65% mAP). This disparity highlights how standard aggregate losses can fail to ensure equitable performance for rare or difficult categories.
Our results demonstrate that FairPO variants, particularly FairPO-CPO, effectively address this challenge by navigating the fairness-performance trade-off in a controlled manner. On MS-COCO, FairPO-CPO achieves a substantial +3.44% mAP gain for $\mathcal{P}$ labels, while performance on $\bar{\mathcal{P}}$ labels dips by a negligible 0.31%. Similarly, on NUS-WIDE, it yields a +3.59% mAP gain for a minor 0.41% drop. This is not an uncontrolled side effect but a managed outcome, enforced by the GRPO mechanism and the constrained loss (Eq. 17), which together prevent significant harm to the non-privileged group while targeting improvements for the privileged one. By directly optimizing the model to rank true labels higher than their specific, dynamically identified confusing counterparts, FairPO sharpens discriminative power for the most challenging cases.
Among the variants, the reference-free FairPO-CPO proves most effective. This directly addresses a potential limitation of DPO-style methods, which can be sensitive to the quality of the reference model. FairPO-CPO’s strong performance, likely due to its dual objective of optimizing both relative preference and absolute correctness (via its NLL-like regularizer), demonstrates its robustness and versatility, making it a more practical choice than the reference-dependent FairPO-DPO or the fixed-margin FairPO-SimPO.
@@ 6 Ablation Studies
Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 55.20 37.15 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (-0.47) (-1.08) (+0.05) (-10.23) (+1.92)
Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT (86.32). Parentheses show change vs Full FairPO-CPO.
Preference Detail Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 (Conf. Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 22.54 35.10 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) (-42.89) (-0.13) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 60.75 34.90 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07) (-4.68) (-0.33)
We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions (Tables 3 and 4). Our analysis of core components (Table 3) confirms the criticality of each part. Replacing the preference loss with BCE reduces privileged gains, removing the $\bar{\mathcal{P}}$ constraint harms non-privileged performance, and using fixed weights instead of GRPO degrades both groups. To further isolate the value of our fairness framework, we tested a Global CPO variant that applies the preference objective to all labels without the privileged/non-privileged split or GRPO. While this proves the preference objective is a strong general-purpose loss (achieving a +2.23% mAP gain over BCE), the full FairPO framework is decisively superior for the targeted fairness task. By applying CPO globally, the performance gains are unfocused, resulting in a 1.21 point lower mAP on the privileged set. This demonstrates that the full FairPO architecture is crucial for focusing the model’s capacity to resolve the hardest discrimination challenges within the privileged group.
Furthermore, the design of the preference objective itself is crucial (Table 4). Restricting it to only Confusing Negatives (for $y_{il}=+1$) causes a profound performance drop, as this neglects the far more common scenario of ranking true negatives below confusing positives for rare labels, making the learning signal exceptionally sparse. Similarly, removing the BCE Fallback for non-confusing instances degrades performance, proving that a standard classification signal on easier cases is vital for model stability. Both aspects of our full formulation are thus essential, demonstrating that while targeted preference optimization is powerful, it must be complemented by standard losses and robust balancing for stable, effective training.
@@ 7 Related Work
Recent efforts in fair MLC address complex challenges like label imbalance impacting tail labels (Guo and Wang, 2021), subjective fairness (Liu et al., 2023), and class-incremental learning (Dong et al., 2025). FairPO contributes a novel approach by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets. It applies distinct, fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and manages the trade-off with a robustness framework, differentiating our targeted approach from prior work.
We adapt recent advances in preference optimization, originally developed for aligning LLMs (Ouyang and Others, 2022; Christiano et al., 2023). Techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and its reference-free variants CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies from preference pairs. Rather than ranking entire outputs, we repurpose these methods to specifically differentiate true label scores from their dynamically identified confusing counterparts within the privileged set, thereby sharpening critical decision boundaries. To balance our objectives, we employ a Group Robust Optimization strategy inspired by Group DRO (Sagawa et al., 2020; Rice et al., 2021) and GRPO (Ramesh et al., 2024). While these methods typically balance performance across data or preference groups, FairPO uniquely defines its groups by our label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$). It then uses GRPO’s adaptive weighting to balance their distinct, custom-formulated loss objectives, providing a principled mechanism for managing the specific fairness-performance trade-offs in our MLC context.

### v4

@@ 2.3 Optimization Algorithm
We solve the minimax objective in Eq. 17 iteratively using alternating mirror descent (Algorithm 1). A key component for stability is loss scaling during the update of the group weights $\alpha$. Since the group losses, $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$, can have different scales and variances, using raw loss values for the mirror ascent step is unstable. To mitigate this, we update $\alpha$ based on the relative change of each group’s loss from its running average, $\bar{\mathcal{L}}_{g}^{\text{avg}}$, ensuring a more stable optimization. Further details are in Appendix E.
@@
6:     Compute privileged loss $\mathcal{L}_{\mathcal{P}}^{(s)}$ for $(x_{i},r)$ (DPO Eq. 13, or BCE Eq. 11).
@@
13:Full details are in Algorithm 2 (see Appendix E).
@@ 3 Experimental Setup
We evaluate FairPO on two multi-label benchmarks: MS-COCO 2014 (lin2015microsoftcococommonobjects) and NUS-WIDE (chua2009nus). To address class imbalance, we partition labels by frequency: the 20) and the remaining 80). While this partitioning is effective, the FairPO framework is general and could use other criteria like domain importance. We assess performance using mAP, Sample F1, and Accuracy, focusing on the mAP gain on the privileged set, , relative to the BCE-SFT baseline. We compare against baselines including Group DRO + BCE (sagawa2019distributionally) and Focal Loss (lin2018focallossdenseobject). Our base model is a frozen Vision Transformer (ViT) (dosovitskiy2021imageworth16x16words) with separate non-linear MLP heads per label, trained with AdamW (loshchilov2019decoupledweightdecayregularization). The trained BCE-SFT baseline provides the reference model parameters. We test three framework variants: FairPO-DPO, FairPO-SimPO, and FairPO-CPO. Full details are in Appendix F.
@@ 4 Results and Analysis
As shown in Tables 1 and 2, FairPO-CPO consistently achieves the highest performance on the privileged group ($\mathcal{P}$), confirming its effectiveness. On MS-COCO, it delivers a statistically significant gain of +3.44% $\Delta\text{mAP}(\mathcal{P})$ over strong baselines like BCE-SFT. Crucially, this targeted improvement comes with only a minor and statistically insignificant performance drop on the non-privileged group ($\bar{\mathcal{P}}$), demonstrating FairPO’s ability to reallocate model capacity without causing measurable harm. The superiority of the CPO variant stems from its robust design: unlike DPO, it is reference-free, and its integrated BCE regularizer provides a crucial signal for absolute correctness that SimPO lacks. This dual objective of learning both relative preferences and absolute scores yields a more stable and effective training process, making FairPO-CPO the most practical choice.
[…]
@@
Method mAP Sample F1 Accuracy $\Delta$ mAP ($\mathcal{P}$) $\Delta$ mAP ($\bar{\mathcal{P}}$) $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ (vs. Focal) (vs. BCE) BCE-SFT (bce_loss) 86.32±0.11 90.65±0.08 61.43±0.15 64.89±0.12 94.89±0.09 98.12±0.05 -2.03 Ref GDRO + BCE (sagawa2019distributionally) 87.92±0.12 90.41±0.09 62.31±0.20 64.75±0.13 95.72±0.10 98.05±0.05 -0.43 -0.24 Focal Loss (lin2018focallossdenseobject) 88.35±0.14 89.81±0.12 63.15±0.16 64.18±0.15 96.11±0.12 97.90±0.07 Ref -0.84 FairPO-DPO 88.02±0.15 89.97±0.11 63.45±0.17 63.65±0.16 97.89±0.13 98.04±0.06 -0.33 -0.68 FairPO-SimPO 87.67±0.18 88.76±0.21 62.34±0.22 63.12±0.19 95.69±0.15 97.78±0.09 -0.68 -1.89 FairPO-CPO 89.76±0.09 90.34±0.07 64.01±0.13 64.32±10 98.03±0.07 98.06±0.05 +1.41 -0.31
Table 2: Performance comparison on NUS-WIDE. Notations are similar to Table 1.
@@
Method mAP Sample F1 Accuracy $\Delta$ mAP ($\mathcal{P}$) $\Delta$ mAP ($\bar{\mathcal{P}}$) $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ (vs. Focal) (vs. BCE) BCE SFT (bce_loss) 63.53±0.15 70.24±0.11 48.12±0.21 55.83±0.14 91.51±0.12 95.22±0.08 -2.28 Ref GDRO + BCE (sagawa2019distributionally) 64.84±0.16 69.91±0.12 49.13±0.22 55.62±0.15 92.11±0.13 95.13±0.09 -0.97 -0.33 Focal Loss (lin2018focallossdenseobject) 65.81±0.17 68.95±0.15 50.33±0.20 54.31±0.18 93.05±0.11 94.75±0.11 Ref -1.29 FairPO-DPO 66.34±0.20 69.05±0.14 51.71±0.25 54.52±0.19 93.92±0.16 95.04±0.09 +0.53 -1.19 FairPO-SimPO 64.11±0.22 68.03±0.24 48.82±0.28 53.81±0.22 91.94±0.18 94.52±0.13 -1.70 -2.21 FairPO-CPO 67.12±0.14 69.83±0.10 52.21±0.19 55.24±0.13 94.31±0.10 95.12±0.08 +1.31 -0.41
@@ 5 Ablation Studies
Ablation studies on MS-COCO with FairPO-CPO confirm that each component is critical (Tables 3 & 4). Removing the preference loss, the non-privileged constraint, or the adaptive GRPO balancing all lead to significant performance degradation. A non-targeted Global CPO variant, while better than standard BCE, is substantially less effective than the full FairPO model, highlighting the necessity of our targeted approach. The design of the preference objective itself is also vital: restricting the preference signal to only Confusing Negatives or removing the BCE Fallback for non-confusing cases both degrade performance and stability. These results demonstrate that while preference optimization is powerful, its effectiveness hinges on the complete, balanced FairPO framework.
@@
Method Variant mAP Sample F1 Accuracy $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (+0.47) (-1.08) (+0.05)
Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT. Parentheses show change vs Full FairPO-CPO.
Preference Detail Variant mAP Sample F1 Accuracy $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 +3.44 (Conf. Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07)
@@ 6 Related Work
Our work is situated at the intersection of three research areas, which we detail further in Appendix C. First, while recent efforts in fair MLC have addressed challenges such as tail labels (wu2023longtailed) and subjective fairness (gao2023learning), FairPO contributes a novel approach by partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets and applying targeted, distinct objectives to each. Second, we adapt modern preference optimization techniques, originally developed for aligning LLMs like DPO (rafailov2023direct), CPO (xu2024contrastivepreferenceoptimizationpushing), and SimPO (meng2024simposimplepreferenceoptimization). Instead of ranking entire outputs, we repurpose these methods to resolve ambiguities between true labels and their dynamically identified confusing counterparts. Finally, to manage the trade-off between our objectives, we employ Group Robust Optimization. Inspired by Group DRO (sagawa2020distributionallyrobustneuralnetworks) and GRPO (ramesh2024group), we uniquely define the groups by our label partitions ($\mathcal{P}$ and $\bar{\mathcal{P}}$) and use GRPO’s adaptive weighting to balance their custom-formulated losses.